\begin{thebibliography}{}

\bibitem[\protect\citename{Agic and Vulic}2019]{agic-vulic-2019-jw300}
Zeljko Agic and Ivan Vulic.
\newblock 2019.
\newblock Jw300: A wide-coverage parallel corpus for low-resource languages.
\newblock In {\em Proceedings of the 57th Annual Meeting of the Association for
  Computational Linguistics}, pages 3204--3210, Florence, Italy, jul.
  Association for Computational Linguistics.

\bibitem[\protect\citename{Bahdanau and Bengio}2015]{attention}
Cho~K. Bahdanau, D. and Y.~Bengio.
\newblock 2015.
\newblock {\em Neural machine translation by jointly learning to align and
  translate}.

\bibitem[\protect\citename{Bender and Friedman}2018]{bender-friedman-2018-data}
Emily~M. Bender and Batya Friedman.
\newblock 2018.
\newblock Data statements for natural language processing: Toward mitigating
  system bias and enabling better science.
\newblock {\em Transactions of the Association for Computational Linguistics},
  6:587--604.

\bibitem[\protect\citename{Brownlee}2017]{brownlee}
Jason Brownlee.
\newblock 2017.
\newblock {\em Deep Learning for Natural Language Processing}.
\newblock Machine Learning Mastery.

\bibitem[\protect\citename{Callison-Burch \bgroup et al.\egroup
  }2006]{callison-burch-etal-2006-evaluating}
Chris Callison-Burch, Miles Osborne, and Philipp Koehn.
\newblock 2006.
\newblock Re-evaluating the role of {B}leu in machine translation research.
\newblock In {\em 11th Conference of the {E}uropean Chapter of the Association
  for Computational Linguistics}, Trento, Italy, April. Association for
  Computational Linguistics.

\bibitem[\protect\citename{Dossou and Emezue}2020]{ffr}
Bonaventure F.~P. Dossou and Chris~C. Emezue.
\newblock 2020.
\newblock Ffr v1.0: Fon-french neural machine translation.

\bibitem[\protect\citename{Fadaee \bgroup et al.\egroup
  }2017]{dataaugmentation}
Marzieh Fadaee, Arianna Bisazza, and Christof Monz.
\newblock 2017.
\newblock Data augmentation for low-resource neural machine translation.

\bibitem[\protect\citename{Graham \bgroup et al.\egroup }2016]{graham}
Yvette Graham, TIMOTHY BALDWIN, Alistair Moffat, and Justin Zobel.
\newblock 2016.
\newblock Can machine translation systems be evaluated by the crowd alone.
\newblock {\em Natural Language Engineering}, pages 1--28, 09.

\bibitem[\protect\citename{Hochreiter and Schmidhuber}1997]{lstm}
Sepp Hochreiter and Jurgen Schmidhuber.
\newblock 1997.
\newblock Long short-term memory.
\newblock {\em Neural Comput.}, 9(8):1735–1780, Nov.

\bibitem[\protect\citename{Koehn and Monz}2006]{koehn}
Philipp Koehn and Christof Monz.
\newblock 2006.
\newblock Manual and automatic evaluation of machine translation between
  european languages.
\newblock In {\em Proceedings of the Workshop on Statistical Machine
  Translation}, StatMT ’06, page 102–121, USA. Association for
  Computational Linguistics.

\bibitem[\protect\citename{Lamba}2020]{lamba}
H.~Lamba.
\newblock 2020.
\newblock Intuitive understanding of attention mechanism in deep learning.
\newblock {\em Medium}.

\bibitem[\protect\citename{Lo \bgroup et al.\egroup
  }2013]{lo-etal-2013-improving}
Chi-kiu Lo, Karteek Addanki, Markus Saers, and Dekai Wu.
\newblock 2013.
\newblock Improving machine translation by training against an automatic
  semantic frame based evaluation metric.
\newblock In {\em Proceedings of the 51st Annual Meeting of the Association for
  Computational Linguistics (Volume 2: Short Papers)}, pages 375--381, Sofia,
  Bulgaria, August. Association for Computational Linguistics.

\bibitem[\protect\citename{Martinus and Abbott}2019]{jade}
Laura Martinus and Jade~Z. Abbott.
\newblock 2019.
\newblock A focus on neural machine translation for african languages.

\bibitem[\protect\citename{NMT}2020]{tensorflow}
Tensorflow NMT.
\newblock 2020.
\newblock Neural machine translation with attention.
\newblock Tensorflow.

\bibitem[\protect\citename{Orife \bgroup et al.\egroup
  }2020a]{Orife2020ImprovingYD}
Iroro Orife, David~Ifeoluwa Adelani, Timi~E. Fasubaa, Victor Williamson,
  Wuraola~Fisayo Oyewusi, Olamilekan Wahab, and Kola Tubosun.
\newblock 2020a.
\newblock Improving yor\`ub\'a diacritic restoration.
\newblock {\em arXiv: Computation and Language}.

\bibitem[\protect\citename{Orife \bgroup et al.\egroup }2020b]{masakhane}
Iroro Orife, Julia Kreutzer, Blessing Sibanda, Daniel Whitenack, Kathleen
  Siminyu, Laura Martinus, Jamiil~Toure Ali, Jade Abbott, Vukosi Marivate,
  Salomon Kabongo, Musie Meressa, Espoir Murhabazi, Orevaoghene Ahia, Elan van
  Biljon, Arshath Ramkilowan, Adewale Akinfaderin, Alp Öktem, Wole Akin,
  Ghollah Kioko, Kevin Degila, Herman Kamper, Bonaventure Dossou, Chris Emezue,
  Kelechi Ogueji, and Abdallah Bashir.
\newblock 2020b.
\newblock Masakhane -- machine translation for africa.

\bibitem[\protect\citename{Papineni \bgroup et al.\egroup
  }2002]{papineni-etal-2002-bleu}
Kishore Papineni, Salim Roukos, Todd Ward, and Wei-Jing Zhu.
\newblock 2002.
\newblock {B}leu: a method for automatic evaluation of machine translation.
\newblock In {\em Proceedings of the 40th Annual Meeting of the Association for
  Computational Linguistics}, pages 311--318, Philadelphia, Pennsylvania, USA,
  July. Association for Computational Linguistics.

\bibitem[\protect\citename{Sennrich \bgroup et al.\egroup
  }2015]{DBLP:journals/corr/SennrichHB15}
Rico Sennrich, Barry Haddow, and Alexandra Birch.
\newblock 2015.
\newblock Neural machine translation of rare words with subword units.
\newblock {\em CoRR}, abs/1508.07909.

\bibitem[\protect\citename{Sutskever \bgroup et al.\egroup }2014]{NIPS20145346}
Ilya Sutskever, Oriol Vinyals, and Quoc~V Le.
\newblock 2014.
\newblock Sequence to sequence learning with neural networks.
\newblock In Z.~Ghahramani, M.~Welling, C.~Cortes, N.~D. Lawrence, and K.~Q.
  Weinberger, editors, {\em Advances in Neural Information Processing Systems
  27}, pages 3104--3112. Curran Associates, Inc.

\bibitem[\protect\citename{Turian \bgroup et al.\egroup
  }2003]{Turian03evaluationof}
Joseph Turian, Luke Shen, and I.~Dan Melamed.
\newblock 2003.
\newblock Evaluation of machine translation and its evaluation.
\newblock In {\em In Proceedings of MT Summit IX}, pages 386--393.

\bibitem[\protect\citename{Vaswani \bgroup et al.\egroup
  }2017]{Vaswani2017AttentionIA}
Ashish Vaswani, Noam Shazeer, Niki Parmar, Jakob Uszkoreit, Llion Jones,
  Aidan~N. Gomez, Lukasz Kaiser, and Illia Polosukhin.
\newblock 2017.
\newblock Attention is all you need.
\newblock {\em ArXiv}, abs/1706.03762.

\bibitem[\protect\citename{Wu \bgroup et al.\egroup }2016]{gleu}
Yonghui Wu, Mike Schuster, Zhifeng Chen, Quoc~V. Le, Mohammad Norouzi, Wolfgang
  Macherey, Maxim Krikun, Yuan Cao, Qin Gao, Klaus Macherey, Jeff Klingner,
  Apurva Shah, Melvin Johnson, Xiaobing Liu, Łukasz Kaiser, Stephan Gouws,
  Yoshikiyo Kato, Taku Kudo, Hideto Kazawa, Keith Stevens, George Kurian,
  Nishant Patil, Wei Wang, Cliff Young, Jason Smith, Jason Riesa, Alex Rudnick,
  Oriol Vinyals, Greg Corrado, Macduff Hughes, and Jeffrey Dean.
\newblock 2016.
\newblock Google's neural machine translation system: Bridging the gap between
  human and machine translation.
\newblock {\em CoRR}, abs/1609.08144.

\bibitem[\protect\citename{Xia \bgroup et al.\egroup }2019]{dataaugmentation1}
Mengzhou Xia, Xiang Kong, Antonios Anastasopoulos, and Graham Neubig.
\newblock 2019.
\newblock Generalized data augmentation for low-resource translation.

\end{thebibliography}

\documentclass[11pt]{article}
\usepackage{coling2018}
\usepackage{times}
\usepackage{url}
\usepackage{latexsym}
\usepackage{tabularx}
\usepackage{multirow}
\usepackage{graphicx}
\usepackage{xcolor}
\usepackage{textcomp}
\usepackage[T4,T1]{fontenc}
\newenvironment{tfour}{\fontencoding{T4}\selectfont}{}
\usepackage{newunicodechar}
\usepackage{enumitem}

\newunicodechar{ɔ}{\fon}

\title{FFR v1.1: Fon-French Neural Machine Translation}
\author{Bonaventure F. P. Dossou \\
 Kazan Federal University\\
\texttt{femipancrace.dossou@gmail.com} \\\And
Chris C. Emezue \\
Kazan Federal University\\
\texttt{chris.emezue@gmail.com} \\}

\begin{document}
\maketitle
\begin{abstract}
All over the world and especially in Africa, researchers are putting efforts into building Neural Machine Translation (NMT) systems to help tackle the language barriers in Africa, a continent of over 2000 different languages. However, the low-resourceness, diacritical, and tonal complexities of African languages are major issues being faced. The FFR project is a major step towards creating a robust translation model from Fon, a very low-resource and tonal language, to French, for research and public use. In this paper, we introduce FFR Dataset, a corpus of Fon-to-French translations, describe the diacritical encoding process, and introduce our FFR v1.1 model, trained on the dataset. The dataset and model are made publicly available at \url{https://github.com/bonaventuredossou/ffr-v1}, to promote collaboration and reproducibility.%and are still being improved daily.
\end{abstract}
\section{The FFR Dataset:  FFR1 and FFR2}
\label{ffrdatasets}
The FFR Dataset is a project to compile a large, growing corpus of carefully cleaned of Fon - French (FFR) parallel sentences for machine translation, and other NLP research-related, projects \cite{ffr}. There are currently two versions of the FFR dataset: the initial FFR dataset (FFR1) and the latest version (FFR2).

The major sources for the creation of FFR1 were JW300 \cite{agic-vulic-2019-jw300} and BeninLangues\footnote{https://beninlangues.com/} with 27980 and 89049 aligned sentences respectively, giving a total of 117,029 parallel sentences. JW300 (JW) contains translations of Jehovah Witness sermons in over 100 languages, while BeninLangues (BL) contains vocabulary words, short expressions, small sentences, complex sentences, proverbs, as well as books of the Bible (Genesis 1 - Psalm 79). 

The initial samples contained various grammatical errors, incorrect and incomplete translations, which were disregarded by standard, rule-based cleaning techniques\footnote{Using Python Regex and String packages (https://docs.python.org/3/library/re.html) and NLTK preprocessing library (https://www.nltk.org/)}. FFR2, obtained after re-evaluation of translations in FFR1 by FFR natives, reduces JW and BL original samples respectively to 26510 and 27465 Fon-French parallel sentences. We also created a data statement \cite{bender-friedman-2018-data} for FFR2, which serves to help give a thorough overview of the dataset. Our data statement can be accessed at \url{https://github.com/bonaventuredossou/ffr-v1/blob/master/FFR-Dataset/Data_Statement_FFR_Dataset.pdf}. The tabular analyses shown in Table \ref{ffrdataset} below serve to give an idea of the range of word lengths for the sentences in FFR1 and FFR2. The maximum number of words-per-sentence for the Fon sentences, $max-fon$, is 109, for FFR1, and 88, for FFR2. That of the French sentences, $max-fr$, is 111 for FFR1 and 76 for FFR2. Therefore, the dataset (both FFR1 and FFR2) has a good range of short, medium and long sentences.
\begin{table}[h]
\caption{ \bf Analysis of Sentences in FFR1}
\label{ffrdataset}
\begin{center}

\begin{tabular}{lllll}
\multicolumn{1}{l}{} &\multicolumn{2}{l}{\bf FFR1} & \multicolumn{2}{l}{\bf FFR2} \\
\hline
\multicolumn{1}{l}{} &\multicolumn{1}{l}{\bf FON} & \multicolumn{1}{l}{\bf FRENCH}&\multicolumn{1}{l}{\bf FON} & \multicolumn{1}{l}{\bf FRENCH} \\
\# Very Short sentences [1-5 words]      &64301          &64255&27470          &30817        \\
\# Short sentences [6-10 words]            &13848                    &17183  &6898                    &12500  \\
\# Medium sentences [11-30 words]             &29113    &29857 &17529    &10582 \\
\# Long sentences [31-($max-fon$ or $max-fr$ )]          &9767           &5734 &2078           &76 \\
Total & \bf 117029 && \bf 53975&

\end{tabular}
\end{center}
\end{table}

%\begin{table}[h]
%\caption{ \bf Analysis of Sentences in FFR2}
%\label{cffrdataset}
%\begin{center}
%
%\begin{tabular}{lll}
%\multicolumn{1}{l}{} &\multicolumn{1}{l}{\bf FON} & %\multicolumn{1}{l}{\bf FRENCH} \\
%\# Very Short sentences [1-5 words]      &27470          %&30817        \\
% \# Short sentences [6-10 words]            &6898        %            &12500   \\
%\# Medium sentences [11-30 words]             &17529    %&10582 \\
%\# Long sentences [31-($max-fon$ or $max-fr$ )]          %&2078           &76 \\
%\end{tabular}
%\end{center}
%\end{table}

%\begin{table}[h]
%\caption{\bf Table showing how the different diacritics affect %the meaning of the Fon word \textit{to}.}
%\label{fonto}
%\begin{center}
%\begin{tabular}{ll}
%\multicolumn{1}{l}{\bf Words with accents} & %\multicolumn{1}{l}{\bf Meaning} \\
%tó & ears \\
%tò & sea  \\
%tô & country \\
%{t\begin{tfour}\m{o}\end{tfour}}&father \\
%\end{tabular}
%\end{center}
%\end{table}
\section{Data Preprocessing}
\label{dateprocess}
Initial analysis of Fon sentences revealed that different accents (or diacritics or tone marking)\footnote{\url{https://en.wikipedia.org/wiki/Fon_language#Tone_marking}} on same words affected their meanings, making it necessary to keep the accents (diacritics) of Fon tokens (words, characters). The importance of encoding diacritics (Diacritical Encoding (DE)) of African languages to NMT has been highlighted by researchers \cite{Orife2020ImprovingYD}, who in their experiments affirmed that DE reduces lexical disambiguation, and helps provide more morphological information to the model. DE was performed using the Normalization Form Canonical Composition (NFC) instead of the Normalization Form Canonical Decomposition (NFD) \footnote{\url{https://unicode.org/reports/tr15/#Norm_Forms}}. With NFC, characters are decomposed and then recomposed by canonical equivalence, while with NFD, they are simply decomposed by canonical equivalence, which removes all accents of Fon tokens. For example, considering the Fon word, \textbf{to}, with its different diacritical meanings, [(t$\acute{o}$,ears),(t$\grave{o}$,sea), (t$\hat{o}$, country), ({t\begin{tfour}\m{o}\end{tfour}},father)], we see that using NFC keeps the diacritics and consequently the meaning of the words, while using NFD, simply gives the word \textit{to} leading to ambiguities in the translation. 

%Considering the sentence \colorbox{pink}{xota ðò akpà {\begin{tfour}\m{o}\end{tfour}} jí é etɛwu mǐ ðó n{\begin{tfour}\m{o}\end{tfour}} ð{\begin{tfour}\m{o}\end{tfour}} nugbǒ}, the NFC result is \colorbox{pink}{xota ðò akpà {\begin{tfour}\m{o}\end{tfour}} jí é etɛwu mǐ ðó n{\begin{tfour}\m{o}\end{tfour}} ð{\begin{tfour}\m{o}\end{tfour}} nugbǒ} i.e. the exact sentence while NFD will produce \colorbox{pink}{xota ðo akpa {\begin{tfour}\m{o}\end{tfour}} ji e etɛwu mi ðo n{\begin{tfour}\m{o}\end{tfour}} ð{\begin{tfour}\m{o}\end{tfour}} nugbo}.%

\section{FFR v1.1 Model Structure and Training}
For our experiments, we used FFR2, described in section \ref{ffrdatasets}, which is an improvement of FFR1. We derived 43719, 4858 and 5398 training, validation and testing samples accordingly. We used the Tensorflow TextTokenizer\footnote{\url{https://www.tensorflow.org/api_docs/python/tf/keras/preprocessing/text}} with \textit{none} filter to tokenize FFR sentences and build the vocabularies (for Fon and French), from which numerical sequences or representations of each FFR sentence pair are built with the Tensorflow Preprocessing package\footnote{\url{https://www.tensorflow.org/api_docs/python/tf/keras/preprocessing/sequence}}, and used to train the model.

The FFR v1.1  model, like the FFR v1.0 \cite{ffr}, is based on the encoder-decoder configuration \cite{NIPS20145346,brownlee,tensorflow}. The encoders and decoders are made up of 128-dimensional gated rectified units (GRUs) recurrent layers \cite{lstm}, with a word embedding layer of dimension 512. A 30-dimensional attention model \cite{NIPS20145346,attention,lamba} was also applied in order to help the model make contextual and correct translations. The code for the model has been open-sourced at \url{https://github.com/bonaventuredossou/ffr-v1/blob/master/model_train_test/fon_fr.py}, to promote reproducibility and similar recent initiaves on machine translation of African languages like \cite{jade,masakhane}. FFR v1.1 model was trained using the Tensorflow v1.14 package \cite{tensorflow}.

\section{Initial Results and Findings}
We evaluated the FFR v1.1 model performance using BLEU \cite{papineni-etal-2002-bleu}, and GLEU \cite{gleu} metrics. GLEU, is a sentence-level evaluation metric similar to BLEU. 
\begin{table}[h]
\caption{Evlauation scores on test data}
\label{testscore}
    \centering
\begin{tabular}{lllll}

%\caption{BLEU and GLEU Scores on Test Data}
\multicolumn{1}{l}{} &\multicolumn{2}{l}{\bf FFR1 } & \multicolumn{2}{l}{\bf FFR2} \\
\hline

\multicolumn{1}{l}{} &\multicolumn{1}{l}{\bf BLEU} & \multicolumn{1}{l}{\bf GLEU} &\multicolumn{1}{l}{\bf BLEU} & \multicolumn{1}{l}{\bf GLEU}\\
Without DE&24.53&13.0&27.80&17.05  \\
With DE&\bf{30.55}&\bf{18.18} &\bf{ 37.15}&\bf{20.85} \\

\end{tabular}
\end{table}
As shown on Table \ref{testscore}, the FFR model, trained on both FFR1 and FFR2 showed an improvement when trained with DE. %The FFR v1.1 model performed better on FFR2 than on FFR1, and this testifies of the effectiveness of the deeper cleaning that has been made.

Table~\ref{sample-table} shows translations of interest from the FFR model sources from FFR2, illustrating the difficulty of predicting Fon words which bear different meanings with different accents. While the model predicted well for \#0 and \#1, it misplaced the meanings for \#2 and \#3.  
\begin{table}[h!]
\caption{Sample predictions and scores}
\label{sample-table}
\begin{center}
 \resizebox{\textwidth}{!}{
\begin{tabular}{ |c|c|c|c|c|c|c| } 
 \hline
\bf ID & \bf 0 &\bf 1 &\bf2&\bf3&\bf4&\bf5\\
 \textcolor{blue}{Source} & \textcolor{blue}{y$\acute{i}$ bo wa} & \textcolor{blue}{yi bo wa}  &\textcolor{blue}{h\begin{tfour}\m{o}\end{tfour}n} &\textcolor{blue}{h\begin{tfour}\m{o}\end{tfour}n} & \textcolor{blue}{s\'a amas\'in d\u o w\u u } &\textcolor{blue}{gb\begin{tfour}\m{e}\end{tfour}}
 \\ 
 \textcolor{brown}{Target} &  \textcolor{brown}{prends et viens} &  \textcolor{brown}{va et viens} & \textcolor{brown}{porte} & \textcolor{brown}{fuire} & \textcolor{brown}{oindre avec un m\'edicament} &\textcolor{brown}{pousser de nouvelles feuilles}  \\
 FFR v1.0 Model & prends et viens & va viens & scorpion &porte &se masser avec le remede &esprit de la vie\\
 BLEU/\bf CMS score & 1.0 & 1.0 & 0.0   &0.0 &0.0/\textbf{0.95} &0.25 /\textbf{ 0.9} \\ 
 \hline
\end{tabular} 
}
\end{center}
\end{table}
\subsection{The Context-Meaning-Similarity (CMS) metric}
Researchers have shown that automatic metrics are not necessarily a good substitute for human assessments of translation quality \cite{Turian03evaluationof,callison-burch-etal-2006-evaluating,graham}, due to issues like lexical-vs-semantic similarity and existence of many possible valid translations for each source sentence \cite{koehn,lo-etal-2013-improving,graham}.
During our experiments, we discovered that the FFR v1.1 model was able to provide predictions that were, although different from the target, similar in context to the target, as seen in sentence \#4. Both \textcolor{brown}{oindre avec un m\'edicament} and \textcolor{brown}{se masser avec le remede} convey the same idea in the context of the source sentence, \textcolor{blue}{s\'a amas\'in d\u o w\u u }.\newline
%Most African languages, due to the , have many dialects Also, there is the problem of bias in human assessment of translations in some African languages, due to the multi-ethnicity and varied dialects of the same language.

This led us to experiment a method we call CMS metric:
\begin{footnotesize}
\begin{enumerate}[noitemsep]
\item A subset of the testing data, consisting of 100 specially selected source, target and predicted sentences, was sent to five FFR natives.
\item \label{tscore} They were first given the source and prediction sentences and asked to give a score, $t \in [0,1]$, on how similar the source and prediction sentences were contextually. Note that this scoring was done with no knowledge of the reference, but through the innate experience of the native speakers.
\item  Then they were given the source and prediction along with the reference sentences and, simillar to step \ref{tscore} above, were instructed to give a score $t_{r}$.
\item Using a parameter, $\alpha$, we calculated the total score $t_{total} = \alpha*t + (1-\alpha)*t_{r}$. This parameter controls the tradeoff between the review of the prediction, when viewed on its own, and that of the prediction when viewed in contextual comaprison to the reference sentence. For our experiment, we set $\alpha = 0.7$, putting more weight on the prediction without the reference comparison.
\item The average of these scores was taken as the CMS score for each of the model\textquotesingle s predictions as given in sentence \#4 in Table~\ref{sample-table}. 
\end{enumerate}
\end{footnotesize}
An interesting feature of the CMS metric is the tradeoff, $\alpha$, which is especially useful for translation assessments in languages that have many dialects (like most African languages) and expressions with various possible contexts (like Fon). 
\section{Conclusion, Future Work and Acknowledgements}
In this paper, we introduced the creation of the FFR dataset: a corpus of Fon-French parallel sentences. We further trained an NMT system, and evaluated the translation quality using both the BLEU metric and our proposed CMS metric. Our project is at the pilot stage and therefore, there is headroom to be explored with the tuning of different architectures, learning schemes, transfer learning, tokenization methods for the FFR project (FFR Dataset, FFR model) improvement. Specifically, we are looking into leveraging monolingual data, encoding with subword units \cite{DBLP:journals/corr/SennrichHB15}, exploring data augmentation for low-resource NMT \cite{dataaugmentation,dataaugmentation1}, and training on a state-of-the-art Transformer model \cite{Vaswani2017AttentionIA}. We owe great thanks to Julia Kreutzer, Jade Abott and the Masakhane Community for their mentorship. We would also like to thank the FFR natives for the good translation services provided.

\blfootnote{
    %
    % for review submission
    %
    \hspace{-0.65cm}  % space normally used by the marker
    % Place licence statement here for the camera-ready version. See
    %Section~\ref{licence} of the instructions for preparing a
    %manuscript.
    %
    % % final paper: en-uk version 
    %
    This work is licensed under a Creative Commons 
    Attribution 4.0 International Licence.
    Licence details:
    \url{http://creativecommons.org/licenses/by/4.0/}.
    % 
    % % final paper: en-us version 
    %
    % \hspace{-0.65cm}  % space normally used by the marker
    % This work is licensed under a Creative Commons 
    % Attribution 4.0 International License.
    % License details:
    % \url{http://creativecommons.org/licenses/by/4.0/}.
}

% include your own bib file like this:
%\bibliographystyle{acl}
%\bibliography{coling2018}
\bibliographystyle{acl}
\bibliography{winlp2020}

\end{document}